OpenStudy (anonymous):

Explain how to graph x + 3y = 9 using the x- and y-intercepts method.

OpenStudy (raden):

for the x-intercept, take y=0 x+3y=9 x+3(0)=9 x+0=9 x=9, so x-intercept will be at (9,0) for y-intercept, take x=0 x+3y=9 0+3y=9 3y=9 y=9/3 y=3, so y-intercept will be at (0,3) put of them in grap and joining both points